# Bridged Posterior: Optimization, Profile Likelihood and a New Approach of Generalized Bayes

### v1

@@ 7 Discussion
In our present examples, we have focused on well-behaved loss functions with unique optima, which can be obtained efficiently with high numerical accuracy. Moving beyond this relatively clean setting merits future study as extensions and generalizations of the bridged posterior. Many problems, such as those with non-convex objectives functions, entail losses featuring multiple local optima (Zhang et al., 2020). The solution $z$ returned by the algorithm at convergence may depend on the choice of initialization $\tilde{z}$ in these cases, where our approach has made use of a well-defined solution as input in a hierarchy. One generalization that may be fruitful is to assign a probability distribution over $\tilde{z}$, enabling us to view the optimization procedure as an algorithm mapping to another distribution for $z$. Second, many popular optimization algorithms, including stochastic variants of schemes such as gradient descent as well as early stopping, may produce approximately optimal solutions. In such cases, it may be more appropriate to model $\Pi(z\mid y,\lambda)$ to carry some uncertainty reflecting numerical errors or stopping criteria, in place of the point mass used in our current formulation. It will be interesting to explore the connections between the bridged posterior and Bayesian probabilistic numerical methodology (Cockayne et al., 2019).
As its name suggests, the proposed methodology aims to bridge ideas between the optimization and Bayesian literature. This task is becoming especially important under the popularization of large or heavily parametrized models such as deep neural networks, and the vital goal for uncertainty quantification as a cornerstone of statistical inference. For these complex models, it quickly becomes impractical to view and model all parameters as random. Our solution offers a practical and theoretically sound alternative—to conditionally optimize a subset of parameters, while treating the rest as random—to bridge this gap.
@@ Appendix A Proofs
The proof of Theorem 2 uses a theorem from Miller [2021]. We provide the complete statement of that theorem as following.
@@
Miller [2021, Theorem 5] Let $\Theta\subseteq\mathbb{R}^{D}$. Let $E\subseteq\Theta$ be open (in $\mathbb{R}^{D}$) and bounded. Fix $\theta_{0}\in E$ and let $\pi:\Theta\rightarrow\mathbb{R}$ be a probability density with respect to Lebesgue measure such that $\pi$ is continuous at $\theta_{0}$ and $\pi(\theta_{0})>0$. Let $f_{n}:\Theta\rightarrow\mathbb{R}$ have continuous third derivatives on $E$. Suppose $f_{n}\rightarrow f$ pointwise for some $f:\Theta\rightarrow\mathbb{R},f^{\prime\prime}(\theta_{0})$ is positive definite, and $(f_{n}^{\prime\prime\prime})$ is uniformly bounded on E. If either of the following two assumptions is satisfied:
@@
    $\hat{l}_{n}$ has continuous third derivatives on $E$. Since $l_{n}$ has continuous third derivatives on $E\times\hat{z}_{n}(E)$ and $\hat{z}_{n}$ has continuous third derivative, we have
@@
    $\hat{l}_{n}\to\hat{l}_{*}$ pointwise on $\Theta$ and $\hat{l}_{*}^{\prime\prime}(\lambda_{0})$ is positive definite. The first part is obvious as $l_{n}\to l_{*}$ and $\hat{z}_{n}\to\hat{z}_{*}$. To show $\hat{l}_{*}^{\prime\prime}(\lambda_{0})$ is positive definite, we have
@@
    where the second equation is using $\frac{\partial l_{*}(\lambda_{0},\zeta)}{\partial\zeta}\big{|}_{\zeta=\hat{z}_{*}(\lambda_{0})}=0$ to cancel out the last term.
@@
    $\hat{l}_{n}^{\prime\prime\prime}$ is uniformly bounded on $E$. Since $l_{n}^{\prime\prime\prime}$ and $\hat{z}_{n}^{\prime\prime\prime}$ are uniformly bounded, by the theorem 7 of Miller [2021], $l_{n}^{\prime},\ l_{n}^{\prime\prime},\hat{z}_{n}^{\prime}$ and $\hat{z}_{n}^{\prime\prime}$ are all uniformly bounded. Hence, $\hat{l}_{n}^{\prime\prime\prime}$ is uniformly bounded by the expansion of $\hat{l}_{n}^{\prime\prime\prime}$:
@@
    $\quad+3[\hat{z}_{n}^{\prime}(\lambda)]^{\top}\frac{\partial^{3}l_{n}(\lambda,\zeta)}{\partial\lambda\partial\zeta^{2}}\bigg{|}_{\zeta=\hat{z}_{n}(\lambda)}\hat{z}_{n}^{\prime}(\lambda)+[\hat{z}_{n}^{\prime}(\lambda)]^{\top}\frac{\partial^{3}l_{n}(\lambda,\zeta)}{\partial\zeta^{3}}\bigg{|}_{\zeta=\hat{z}_{n}(\lambda)}[\hat{z}_{n}^{\prime}(\lambda)]^{2}$
    $\quad+3[\hat{z}_{n}^{\prime}(\lambda)]^{\top}\frac{\partial^{2}l_{n}(\lambda,\zeta)}{\partial\zeta^{2}}\bigg{|}_{\zeta=\hat{z}_{n}(\lambda)}\hat{z}_{n}^{\prime\prime}(\lambda)+3\frac{\partial^{2}l_{n}(\lambda,\zeta)}{\partial\lambda\partial\zeta}\bigg{|}_{\zeta=\hat{z}_{n}(\lambda)}\hat{z}_{n}^{\prime\prime}(\lambda)+\frac{\partial l_{n}(\lambda,\zeta)}{\partial\zeta}\bigg{|}_{\zeta=\hat{z}_{n}(\lambda)}\hat{z}_{n}^{\prime\prime\prime}(\lambda).$
@@
    for some compact $K\subseteq E$ with $\lambda_{0}$ in the interior of $K$, $\hat{l}_{*}(\lambda)>\hat{l}_{*}(\lambda_{0})$ for all $\lambda\in K\backslash\{\lambda_{0}\}$ and $\liminf_{n}\inf_{\lambda\in\Theta\backslash K}\hat{l}_{n}(\lambda)>\hat{l}_{*}(\lambda_{0})$.
By Theorem 4 with $f_{n}=-\hat{l}_{n}$ and $f=-\hat{l}_{*}$, 1–4 complete the proof. ∎
Proof of Corollary 1. Now we consider the posterior distribution of CDLV $\hat{z}_{n}(\lambda)$. Since the transformation $\hat{z}_{n}$ is not between two Euclidean spaces with equal dimensions, it is hard to calculate the density function of $\hat{z}_{n}(\lambda)$ when $\lambda\sim\Pi_{n}(\lambda\mid y)$. However, we can use the delta method to find the asymptotic distribution of $\sqrt{n}$-adjusted $\hat{z}_{n}(\lambda)$. Since convergence in total variation implies convergence in distribution (weak convergence), the random vector $\sqrt{n}(\lambda-\lambda_{n})\rightsquigarrow\text{N}(0,H_{0}^{-1})$. Using delta method, we can prove $\sqrt{n}\{\hat{z}_{n}(\lambda)-\hat{z}_{n}(\lambda_{n})\}\rightsquigarrow\text{N}(0,H_{z}^{-1})$ where $H_{z}^{-1}=\hat{z}^{\prime}_{*}(\lambda_{0})H_{0}^{-1}\hat{z}^{\prime}_{*}(\lambda_{0})^{\top}$. When $g_{n}(\zeta,y_{1:n};\lambda)=-L(y_{1:n},\zeta;\lambda)$ with that the bridged posterior coincides with the profile likelihood, the asymptotic variance above can be represented by the second derivatives of $l_{*}$. We now show that $H_{0}=\hat{l}^{\prime\prime}_{*}(\lambda_{0})=l_{*0,\lambda\lambda}-l_{*0,\lambda\zeta}l_{*0,\zeta\zeta}^{-1}l_{*0,\zeta\lambda}$, where $l_{*0,\lambda\lambda},l_{*0,\zeta\zeta},l_{*0,\zeta\lambda},l_{*0,\lambda\zeta}$ are respectively the second partial derivatives $\frac{\partial^{2}l_{*}(\lambda,\zeta)}{\partial\lambda^{2}},\frac{\partial^{2}l_{*}(\lambda,\zeta)}{\partial\zeta^{2}},\frac{\partial^{2}l_{*}(\lambda,\zeta)}{\partial\zeta\partial\lambda},\frac{\partial^{2}l_{*}(\lambda,\zeta)}{\partial\lambda\partial\zeta}$ evaluating at $\lambda=\lambda_{0},\zeta=\hat{z}_{*}(\lambda_{0})$. Since in this case $\frac{\partial l_{n}(\lambda,\zeta)}{\partial\zeta}|_{\zeta=\hat{z}_{n}(\lambda)}=0$, we have
@@
By Miller [2021, Theorem 7], we have $l_{n}^{\prime\prime}\to l_{*}^{\prime\prime}$. Letting $n\to\infty$ and $\lambda=\lambda_{0}$, we have $\hat{z}^{\prime}_{*}(\lambda_{0})=-l_{*0,\zeta\zeta}^{-1}l_{*0,\zeta\lambda}$. Now
@@
Letting $n\to\infty$ and $\lambda=\lambda_{0}$, we have the result $\hat{l}^{\prime\prime}_{*}(\lambda_{0})=l_{*0,\lambda\lambda}-l_{*0,\lambda\zeta}l_{*0,\zeta\zeta}^{-1}l_{*0,\zeta\lambda}$. We assume that both $l_{*0,\lambda\lambda}$ and $l_{*0,\zeta\zeta}$ are positive definite. The asymptotic variance of $\sqrt{n}\{\hat{z}_{n}(\lambda)-\hat{z}_{n}(\lambda_{n})\}$ is thus $\hat{z}^{\prime}_{*}(\lambda_{0})H_{0}^{-1}\hat{z}^{\prime}_{*}(\lambda_{0})^{\top}=l_{*0,\zeta\zeta}^{-1}l_{*0,\zeta\lambda}(l_{*0,\lambda\lambda}-l_{*0,\lambda\zeta}l_{*0,\zeta\zeta}^{-1}l_{*0,\zeta\lambda})^{-1}l_{*0,\lambda\zeta}l_{*0,\zeta\zeta}^{-1}$. If we treat both $\lambda$ and $\zeta$ as interest parameters in the likelihood $L(y,\zeta;\lambda)$ with some joint prior, then Miller [2021] proves $\sqrt{n}\bigl{(}[\lambda\ \zeta]^{\top}-[\lambda_{n}\ \zeta_{n}]^{\top}\bigr{)}\rightsquigarrow N(0,\tilde{H}_{0}^{-1})$ for some sequences $\lambda_{n}$ and $\zeta_{n}$ when $(\lambda,\zeta)\sim\Pi(\lambda,\zeta\mid y)$, where $\tilde{H}_{0}=l^{\prime\prime}_{*}(\lambda_{0},\zeta_{0})$ and $\zeta_{0}$ is a fixed ground-truth of $\zeta$. Marginally, the asymptotic variance of $\sqrt{n}(\zeta-\zeta_{n})$ is the $\zeta$-block of $\tilde{H}_{0}^{-1}$, which is equal to $(l_{*0,\zeta\zeta}-l_{*0,\zeta\lambda}l_{*0,\lambda\lambda}^{-1}l_{*0,\lambda\zeta})^{-1}=l_{*0,\zeta\zeta}^{-1}+l_{*0,\zeta\zeta}^{-1}l_{*0,\zeta\lambda}(l_{*0,\lambda\lambda}-l_{*0,\lambda\zeta}l_{*0,\zeta\zeta}^{-1}l_{*0,\zeta\lambda})^{-1}l_{*0,\lambda\zeta}l_{*0,\zeta\zeta}^{-1}$. Hence, the asymptotic variance of the $j$-th element of $\sqrt{n}(\zeta-\zeta_{n})$ is strictly greater than the one of the $j$-th element of $\sqrt{n}\{\hat{z}_{n}(\lambda)-\hat{z}_{n}(\lambda_{n})\}$.
@@
and define $g_{0}(x)=\exp\{-x^{\top}H_{0}x/2\}\pi_{0}(\lambda_{0})$. We first show that $\int_{\mathbb{R}^{d}}|g_{n}(x)-g_{0}(x)|\,\textup{d}x\xrightarrow[n\to\infty]{P_{\lambda_{0},\zeta_{0}}}0$. Denote the $\epsilon$ chosen from Lemma 1 as $\epsilon_{0}$. Since $\pi_{0}$ is continuous at $\lambda_{0}$, we choose sufficiently small $\epsilon\in(0,\epsilon_{0}/2)$ such that $\pi_{0}(\lambda)\leq 2\pi_{0}(\lambda_{0})$ for all $\lambda\in B_{2\epsilon}(\lambda_{0})$. Let $\delta$ be the number from (12). Murphy and Van der Vaart [2000, Corollary 1] shows that $\sqrt{n}\|\lambda_{n}-\lambda_{0}\|$ is bounded in probability and for all $\lambda\in B_{\epsilon_{0}}(\lambda_{0})$ and large enough $n$ such that $\lambda_{n}\in B_{\epsilon_{0}}(\lambda_{0})$,
@@
When $\|x\|<\epsilon\sqrt{n}$, for $n$ large enough we have $\|(\lambda_{n}+x/\sqrt{n})-\lambda_{0}\|<\|\lambda_{n}-\lambda_{0}\|+\epsilon<2\epsilon$. By the choice of $\epsilon$, we have $\pi_{0}(\lambda_{n}+x/\sqrt{n})\leq 2\pi_{0}(\lambda_{0})$. Since $(\|x\|+1)^{2}\leq 2\|x\|^{2}+2$, we have $o_{P_{\lambda_{0},\zeta_{0}}}(1)(\|x\|+1)^{2}\leq\alpha(\|x\|^{2}+1)/2=\alpha x^{\top}x/2+\alpha/2$. Hence $g_{n}(x)\leq h_{n}(x)$ with probability converge to $1$ for $n$ sufficiently large, combining with (12) when $\|x\|\geq\epsilon\sqrt{n}$. Also, $h_{n}(x)\to h_{0}(x)=\exp\{-x^{\top}A_{0}x/2+\alpha/2\}2\pi_{0}(\lambda_{0})$ pointwise. Now,
@@
The second term is less than $\int_{\mathbb{R}^{d}}e^{-n\delta/2}\pi_{0}(\lambda_{n}+x/\sqrt{n})\,\textup{d}x=e^{-n\delta/2}\int_{\mathbb{R}^{d}}\pi_{0}(\lambda)n^{d/2}\,\textup{d}\lambda=e^{-n\delta/2}n^{d/2}\to 0$, while the first term monotonically converges to $\int_{\mathbb{R}^{d}}h_{0}(x)\,\textup{d}x$. Since $g_{n},g_{0},h_{n},h_{0}$ are integrable, by the generalized dominated convergence theorem (the version for convergence in probability), we have $\int_{\mathbb{R}^{d}}|g_{n}(x)-g_{0}(x)|\,\textup{d}x\xrightarrow[n\to\infty]{P_{\lambda_{0},\zeta_{0}}}0$ and $\int_{\mathbb{R}^{d}}g_{n}(x)\,\textup{d}x\xrightarrow[n\to\infty]{P_{\lambda_{0},\zeta_{0}}}\int_{\mathbb{R}^{d}}g_{0}(x)\,\textup{d}x$. Let $a_{n}=1/\int_{\mathbb{R}^{d}}g_{n}(x)\,\textup{d}x$ and $a_{0}=1/\int_{\mathbb{R}^{d}}g_{0}(x)\,\textup{d}x$. Then $a_{n}\to a_{0}$ in $P_{\lambda_{0},\zeta_{0}}$-probability, and thus
@@
This proves $\text{d}_{\text{TV}}\bigl{\{}q_{n},\mathcal{N}\bigl{(}0,H_{0}^{-1}\bigr{)}\bigr{\}}\xrightarrow[n\to\infty]{P_{\lambda_{0},\zeta_{0}}}0$, and (13) follows from [Miller, 2021, Lemma 28]. ∎

### v2

@@ 7 Discussion
In our present examples, we have focused on well-behaved loss functions with unique optima, which can be obtained efficiently with high numerical accuracy. Moving beyond this relatively clean setting merits future study as extensions and generalizations of the bridged posterior. Many problems, such as those with non-convex objectives functions, entail losses featuring multiple local optima (Zhang et al., 2020). The solution $z$ returned by the algorithm at convergence may depend on the choice of initialization $\tilde{z}$ in these cases, where our approach has made use of a well-defined solution as input in a hierarchy. One generalization that may be fruitful is to assign a probability distribution over $\tilde{z}$, enabling us to view the optimization procedure as an algorithm mapping to another distribution for $z$. Second, many popular optimization algorithms, including stochastic variants of schemes such as gradient descent and early stopping, may produce approximately optimal solutions. In such cases, it may be more appropriate to model $\Pi(z\mid y,\lambda)$ to carry some uncertainty reflecting numerical errors or stopping criteria, in place of the point mass used in our current formulation. It is interesting to explore further connections to areas including Bayesian probabilistic numerical methodology (Cockayne et al., 2019) and optimization-based frequentist confidence intervals for constrained problems (Batlle et al., 2023).
Empirically we observe a faster mixing rate of the Markov chain for the bridged posterior, compared to the chain targeting the joint posterior under canonical full likelihood. A related theory can be found in Liu (1994), who proved that the Markov chain targeting the integrated posterior can have a faster mixing rate than the chain targeting the associated joint posterior. On the other hand, this intuition is not directly extendable to our case, as the bridged posterior and the integrated posterior are often not the same. Therefore, a rigorous analysis on comparing the mixing rates of the two Markov chains remains illusive at this point, and can be pursued as a future work.
@@ Appendix A Proofs
The proof of Theorem 2 uses a theorem from Miller (2021). We provide the complete statement of that theorem as following.
@@
Miller (2021, Theorem 5) Let $\Theta\subseteq\mathbb{R}^{D}$. Let $E\subseteq\Theta$ be open (in $\mathbb{R}^{D}$) and bounded. Fix $\theta_{0}\in E$ and let $\pi:\Theta\rightarrow\mathbb{R}$ be a probability density with respect to Lebesgue measure such that $\pi$ is continuous at $\theta_{0}$ and $\pi(\theta_{0})>0$. Let $f_{n}:\Theta\rightarrow\mathbb{R}$ have continuous third derivatives on $E$. Suppose $f_{n}\rightarrow f$ pointwise for some $f:\Theta\rightarrow\mathbb{R},f^{\prime\prime}(\theta_{0})$ is positive definite, and $(f_{n}^{\prime\prime\prime})$ is uniformly bounded on E. If either of the following two assumptions is satisfied:
@@
    $\hat{l}_{n}$ has continuous third derivatives on $E$. Since $l_{n}$ has continuous third derivatives on $E\times\hat{z}_{n}(E)$ and $\hat{z}_{n}$ has continuous third derivative by Assumption (A1), we have
@@
    $\hat{l}_{n}\to\hat{l}_{*}$ pointwise on $\Theta$ and $-\hat{l}_{*}^{\prime\prime}(\lambda_{0})$ is positive definite. The first part is by Assumption (A2), $l_{n}\to l_{*}$ and $\hat{z}_{n}\to\hat{z}_{*}$. To show $-\hat{l}_{*}^{\prime\prime}(\lambda_{0})$ is positive definite, we have
@@
    where the second equation is using $\frac{\partial l_{*}(\lambda_{0},\zeta)}{\partial\zeta}\big{|}_{\zeta=\hat{z}_{*}(\lambda_{0})}=0$ to cancel out the last term. Note that $-l_{*}^{\prime\prime}(\lambda_{0},\hat{z}_{*}(\lambda_{0}))$ is positive definite by Assumption (A3).
@@
    $\hat{l}_{n}^{\prime\prime\prime}$ is uniformly bounded on $E$. Since $l_{n}^{\prime\prime\prime}$ and $\hat{z}_{n}^{\prime\prime\prime}$ are uniformly bounded, by the theorem 7 of Miller (2021), $l_{n}^{\prime},\ l_{n}^{\prime\prime},\hat{z}_{n}^{\prime}$ and $\hat{z}_{n}^{\prime\prime}$ are all uniformly bounded. Hence, $\hat{l}_{n}^{\prime\prime\prime}$ is uniformly bounded by the expansion of $\hat{l}_{n}^{\prime\prime\prime}$:
@@
    $+3[\hat{z}_{n}^{\prime}(\lambda)]^{\top}\frac{\partial^{3}l_{n}(\lambda,\zeta)}{\partial\lambda\partial\zeta^{2}}\bigg{|}_{\zeta=\hat{z}_{n}(\lambda)}\hat{z}_{n}^{\prime}(\lambda)+[\hat{z}_{n}^{\prime}(\lambda)]^{\top}\frac{\partial^{3}l_{n}(\lambda,\zeta)}{\partial\zeta^{3}}\bigg{|}_{\zeta=\hat{z}_{n}(\lambda)}[\hat{z}_{n}^{\prime}(\lambda)]^{2}$
    $+3[\hat{z}_{n}^{\prime}(\lambda)]^{\top}\frac{\partial^{2}l_{n}(\lambda,\zeta)}{\partial\zeta^{2}}\bigg{|}_{\zeta=\hat{z}_{n}(\lambda)}\hat{z}_{n}^{\prime\prime}(\lambda)+3\frac{\partial^{2}l_{n}(\lambda,\zeta)}{\partial\lambda\partial\zeta}\bigg{|}_{\zeta=\hat{z}_{n}(\lambda)}\hat{z}_{n}^{\prime\prime}(\lambda)$
    $+\frac{\partial l_{n}(\lambda,\zeta)}{\partial\zeta}\bigg{|}_{\zeta=\hat{z}_{n}(\lambda)}\hat{z}_{n}^{\prime\prime\prime}(\lambda).$
@@
    By Assumption (A4), for some compact $K\subseteq E$ with $\lambda_{0}$ in the interior of $K$, $\hat{l}_{*}(\lambda)<\hat{l}_{*}(\lambda_{0})$ for all $\lambda\in K\backslash\{\lambda_{0}\}$ and $\limsup_{n}\sup_{\lambda\in\Theta\backslash K}\hat{l}_{n}(\lambda)<\hat{l}_{*}(\lambda_{0})$.
By Theorem 4 with $f_{n}=-\hat{l}_{n}$ and $f=-\hat{l}_{*}$, the above 1–4 complete the proof. ∎
Proof of Corollary 1. We use the delta method to find the asymptotic distribution of $\sqrt{n}$-adjusted $\hat{z}_{n}(\lambda)$. Since convergence in total variation implies convergence in distribution (weak convergence), the random vector $\sqrt{n}(\lambda-\lambda_{n})\rightsquigarrow\text{N}(0,H_{0}^{-1})$. Using delta method, we can prove $\sqrt{n}\{\hat{z}_{n}(\lambda)-\hat{z}_{n}(\lambda_{n})\}\rightsquigarrow\text{N}(0,H_{z}^{-1})$ where $H_{z}^{-1}=\hat{z}^{\prime}_{*}(\lambda_{0})H_{0}^{-1}\hat{z}^{\prime}_{*}(\lambda_{0})^{\top}$.
When $g_{n}(\zeta,y_{1:n};\lambda)=-L(y_{1:n},\zeta;\lambda)$ with that the bridged posterior coincides with the profile likelihood, the asymptotic variance above can be represented by the second derivatives of $l_{*}$. We first show that $-H_{0}=\hat{l}^{\prime\prime}_{*}(\lambda_{0})=l_{*,\lambda_{0}\lambda_{0}}-l_{*,\lambda_{0}\zeta_{0}}l_{*,\zeta_{0}\zeta_{0}}^{-1}l_{*,\zeta_{0}\lambda_{0}}$, where $l_{*,\lambda_{0}\lambda_{0}},l_{*,\zeta_{0}\zeta_{0}},l_{*,\zeta_{0}\lambda_{0}},l_{*,\lambda_{0}\zeta_{0}}$ are respectively the second partial derivatives ${\partial^{2}l_{*}(\lambda,\zeta)}/{\partial\lambda^{2}}$, ${\partial^{2}l_{*}(\lambda,\zeta)}/{\partial\zeta^{2}}$, ${\partial^{2}l_{*}(\lambda,\zeta)}/{\partial\zeta\partial\lambda}$, ${\partial^{2}l_{*}(\lambda,\zeta)}/{\partial\lambda\partial\zeta}$ evaluating at $\lambda=\lambda_{0},\zeta=\zeta_{0}=\hat{z}_{*}(\lambda_{0})$.
Since in this case $\frac{\partial l_{n}(\lambda,\zeta)}{\partial\zeta}|_{\zeta=\hat{z}_{n}(\lambda)}=0$, we have
@@
By Miller (2021, Theorem 7), we have $l_{n}^{\prime\prime}\to l_{*}^{\prime\prime}$. Letting $n\to\infty$ and $\lambda=\lambda_{0}$, we have $\hat{z}^{\prime}_{*}(\lambda_{0})=-l_{*,\zeta_{0}\zeta_{0}}^{-1}l_{*,\zeta_{0}\lambda_{0}}$. Now
@@
Letting $n\to\infty$ and $\lambda=\lambda_{0}$, we have the result $\hat{l}^{\prime\prime}_{*}(\lambda_{0})=l_{*,\lambda_{0}\lambda_{0}}-l_{*,\lambda_{0}\zeta_{0}}l_{*,\zeta_{0}\zeta_{0}}^{-1}l_{*,\zeta_{0}\lambda_{0}}$.
By Assumption (A3), both $-l_{*,\lambda_{0}\lambda_{0}}$ and $-l_{*,\zeta_{0}\zeta_{0}}$ are positive definite. The asymptotic variance of $\sqrt{n}\{\hat{z}_{n}(\lambda)-\hat{z}_{n}(\lambda_{n})\}$ is thus
$\hat{z}^{\prime}_{*}(\lambda_{0})H_{0}^{-1}\hat{z}^{\prime}_{*}(\lambda_{0})^{\top}=-l_{*,\zeta_{0}\zeta_{0}}^{-1}l_{*,\zeta_{0}\lambda_{0}}(l_{*,\lambda_{0}\lambda_{0}}-l_{*,\lambda_{0}\zeta_{0}}l_{*,\zeta_{0}\zeta_{0}}^{-1}l_{*,\zeta_{0}\lambda_{0}})^{-1}l_{*,\lambda_{0}\zeta_{0}}l_{*,\zeta_{0}\zeta_{0}}^{-1}$.
If we treat the latent variable $\zeta$ as non-deterministic in the likelihood $L(y,\zeta;\lambda)$ with some prior, then Miller (2021) proves $\sqrt{n}\bigl{(}[\lambda\ \zeta]^{\top}-[\lambda_{n}\ \zeta_{n}]^{\top}\bigr{)}\rightsquigarrow N(0,\tilde{H}_{0}^{-1})$ for some sequences $\lambda_{n}$ and $\zeta_{n}$ when $(\lambda,\zeta)\sim\Pi(\lambda,\zeta\mid y)$, where $\tilde{H}_{0}=-l^{\prime\prime}_{*}(\lambda_{0},\zeta_{0})$ and $\zeta_{0}=\hat{z}_{*}(\lambda_{0})$ is a fixed ground-truth of $\zeta$. Marginally, the asymptotic variance of $\sqrt{n}(\zeta-\zeta_{n})$ is the $\zeta$-block of $\tilde{H}_{0}^{-1}$, which is equal to
$-(l_{*,\zeta_{0}\zeta_{0}}-l_{*,\zeta_{0}\lambda_{0}}l_{*,\lambda_{0}\lambda_{0}}^{-1}l_{*,\lambda_{0}\zeta_{0}})^{-1}=-l_{*,\zeta_{0}\zeta_{0}}^{-1}-l_{*,\zeta_{0}\zeta_{0}}^{-1}l_{*,\zeta_{0}\lambda_{0}}(l_{*,\lambda_{0}\lambda_{0}}-l_{*,\lambda_{0}\zeta_{0}}l_{*,\zeta_{0}\zeta_{0}}^{-1}l_{*,\zeta_{0}\lambda_{0}})^{-1}l_{*,\lambda_{0}\zeta_{0}}l_{*,\zeta_{0}\zeta_{0}}^{-1}.$
Since $-l_{*,\zeta_{0}\zeta_{0}}^{-1}$ is positive definite, the asymptotic variance of the $j$-th element of $\sqrt{n}(\zeta-\zeta_{n})$ is strictly greater than the one of the $j$-th element of $\sqrt{n}\{\hat{z}_{n}(\lambda)-\hat{z}_{n}(\lambda_{n})\}$. ∎
@@
and define $g_{0}(x)=\exp\{-x^{\top}H_{0}x/2\}\pi_{0}(\lambda_{0})$. We first show that $\int_{\mathbb{R}^{d}}|g_{n}(x)-g_{0}(x)|\,\textup{d}x\xrightarrow[n\to\infty]{P_{\lambda_{0},\zeta_{0}}}0$. Denote the $\epsilon$ chosen from Lemma 1 as $\epsilon_{0}$. Since $\pi_{0}$ is continuous at $\lambda_{0}$, we choose sufficiently small $\epsilon\in(0,\epsilon_{0}/2)$ such that $\pi_{0}(\lambda)\leq 2\pi_{0}(\lambda_{0})$ for all $\lambda\in B_{2\epsilon}(\lambda_{0})$. Let $\delta$ be the number from the condition.
Murphy and Van der Vaart (2000, Corollary 1) shows that $\sqrt{n}\|\lambda_{n}-\lambda_{0}\|$ is bounded in probability and for all $\lambda\in B_{\epsilon_{0}}(\lambda_{0})$ and large enough $n$ such that $\lambda_{n}\in B_{\epsilon_{0}}(\lambda_{0})$,
@@
When $\|x\|<\epsilon\sqrt{n}$, for $n$ large enough we have $\|(\lambda_{n}+x/\sqrt{n})-\lambda_{0}\|<\|\lambda_{n}-\lambda_{0}\|+\epsilon<2\epsilon$. By the choice of $\epsilon$, we have $\pi_{0}(\lambda_{n}+x/\sqrt{n})\leq 2\pi_{0}(\lambda_{0})$. Since $(\|x\|+1)^{2}\leq 2\|x\|^{2}+2$, we have $o_{P_{\lambda_{0},\zeta_{0}}}(1)(\|x\|+1)^{2}\leq\alpha(\|x\|^{2}+1)/2=\alpha x^{\top}x/2+\alpha/2$. Hence $g_{n}(x)\leq h_{n}(x)$ with probability converge to $1$ for $n$ sufficiently large, combining with the condition in theorem when $\|x\|\geq\epsilon\sqrt{n}$.
Also, $h_{n}(x)\to h_{0}(x)=\exp\{-x^{\top}A_{0}x/2+\alpha/2\}2\pi_{0}(\lambda_{0})$ pointwise. Now,
@@
The second term is less than $\int_{\mathbb{R}^{d}}e^{-n\delta/2}\pi_{0}(\lambda_{n}+x/\sqrt{n})\,\textup{d}x=e^{-n\delta/2}\int_{\mathbb{R}^{d}}\pi_{0}(\lambda)n^{d/2}\,\textup{d}\lambda=e^{-n\delta/2}n^{d/2}\to 0$, while the first term monotonically converges to $\int_{\mathbb{R}^{d}}h_{0}(x)\,\textup{d}x$. Since $g_{n},g_{0},h_{n},h_{0}$ are integrable, by the generalized dominated convergence theorem (the version for convergence in probability), we have $\int_{\mathbb{R}^{d}}|g_{n}(x)-g_{0}(x)|\,\textup{d}x\xrightarrow[n\to\infty]{P_{\lambda_{0},\zeta_{0}}}0$ and $\int_{\mathbb{R}^{d}}g_{n}(x)\,\textup{d}x\xrightarrow[n\to\infty]{P_{\lambda_{0},\zeta_{0}}}\int_{\mathbb{R}^{d}}g_{0}(x)\,\textup{d}x$.
Let $a_{n}=1/\int_{\mathbb{R}^{d}}g_{n}(x)\,\textup{d}x$ and $a_{0}=1/\int_{\mathbb{R}^{d}}g_{0}(x)\,\textup{d}x$. Then $a_{n}\to a_{0}$ in $P_{\lambda_{0},\zeta_{0}}$-probability, and thus
@@
This proves $\text{d}_{\text{TV}}\bigl{\{}q_{n},\mathcal{N}\bigl{(}0,H_{0}^{-1}\bigr{)}\bigr{\}}\xrightarrow[n\to\infty]{P_{\lambda_{0},\zeta_{0}}}0$, and $\int_{B_{\epsilon}(\lambda_{0})}\pi_{n}(\lambda)\,\textup{d}\lambda\xrightarrow[n\to\infty]{P_{\lambda_{0},\zeta_{0}}}1$ follows from (Miller, 2021, Lemma 28). ∎